Shaketa:

Adam's average homework score, worth 40% of his grade, is 76%. His average unit test score, worth another 40%, was 81%. The final makes up the rest of his total grade. What is the minimum number of questions he need to answer correctly on the 50-question final exam to earn a B (80%) in the class?

Angle:

Hello Shaketa! Welcome to QuestionCove :) Give me a minute to read your question

Angle:

ok, so here is the process for your question: We want to work with the decimal representations of the percents ` (percent)*(grade) + (percent)*(grade) + (percent)*(grade) = Final percent ` If I plug in what I know, then we have: ` (0.4)*(76) + (0.4)*(81) + (0.2)*(?) = 80 `

Angle:

I am using 0.4 because that is the decimal version of 40% and I used 0.2 for the other one because the final is worth 20%

Angle:

The ? value will be the percent that Adam needs to get on his final -> convert it to a decimal and multiply by 50 to figure out the number of questions

Angle:

Let me know if you need me to take you through the smaller steps - I just covered the main ideas

Shaketa:

Adam's average homework score, worth 40% of his grade, is 76%. His average unit test score, worth another 40%, was 81%. The final makes up the rest of his total grade. What is the minimum number of questions he need to answer correctly on the 50-question final exam to earn a B (80%) in the class? 40 questions 45 questions 42 questions 43 questions

Angle:

Thanks for including your options if you follow the steps that I outlined, you'll be able to get your answer

Angle:

` (0.4)*(76) + (0.4)*(81) + (0.2)*(?) = 80 ` let's first simplify this math, what would you get?